# Chiral anomaly-induced nonlinear Hall effect in spin-orbit coupled noncentrosymmetric metals

### v2

@@ II.1 Maxwell Boltzmann Transport formalism
where $f^{\chi}_{\mathbf{k}}=f_{0}+g^{1,\chi}_{\mathbf{k}}+g^{2,\chi}_{\mathbf{k}}+O(E^{3})$. $f_{0}$ is the standard Fermi-Dirac distribution, $g^{1,\chi}_{\mathbf{k}}$ and $g^{2,\chi}_{\mathbf{k}}$ are deviation up to the first order and second order in electric field $E$, respectively, which can be written in following form  [41, 55, 56]:
@@
Here, ‘$n$’ is the impurity concentration, $\mathcal{V}$’ is the system volume, $u^{\chi}(\mathbf{k})$ is the spinor wave function, $U^{\chi\chi^{\prime}}_{\mathbf{k}\mathbf{k}^{\prime}}$ is the scattering potential profile, and $\epsilon_{F}$ is the Fermi energy. We choose $U^{\chi\chi^{\prime}}_{\mathbf{k}\mathbf{k}^{\prime}}$ such that it can include both magnetic and non-magnetic point-like scattering centers. In general, $U^{\chi\chi^{\prime}}_{\mathbf{k}\mathbf{k}^{\prime}}=U^{\chi\chi^{\prime}}\sigma_{i}$ with $i=0,x,y,z$, where $U^{\chi\chi^{\prime}}$ distinguishes the interband ($\chi\neq\chi^{\prime}$) and intraband ($\chi=\chi^{\prime}$) scattering. In this work, we consider the geometry represented in Fig. 3(a), where the electric field is fixed along the $z$-direction, and the magnetic field is rotated in the $xz$-plane, forming an angle $\gamma$ with the $x$-axis. Including the topological nature of the Hamiltonian, in the presence of electric ($\mathbf{E}$) and magnetic ($\mathbf{B}$) fields, semiclassical dynamics of the Bloch electrons are modified and governed by the following equation [21, 57]:
@@
where, $\mathbf{v}_{\mathbf{k}}^{\chi}=\frac{1}{\hbar}\frac{\partial\epsilon^{\chi}(\mathbf{k})}{\partial\mathbf{k}}$ is band velocity, $\bm{\Omega}^{\chi}_{\mathbf{k}}=i\nabla_{\mathbf{k}}\times\langle u^{\chi}(\mathbf{k})|\nabla_{\mathbf{k}}|u^{\chi}(\mathbf{k})\rangle$ is the Berry curvature, and $\mathcal{D}^{\chi}=(1+e\mathbf{B}\cdot\bm{\Omega}^{\chi}/\hbar)^{-1}$ is factor by which density of states is modified due to presence of the Berry curvature. Self-rotation of the Bloch wave packet also gives rise to an orbital magnetic moment (OMM) $\mathbf{m}^{\chi}_{\mathbf{k}}=-\frac{ie}{2\hbar}\text{Im}\langle\nabla_{\mathbf{k}}u^{\chi}|[\epsilon_{0}(\mathbf{k})-\hat{H}^{\chi}(\mathbf{k})]|\nabla_{\mathbf{k}}u^{\chi}\rangle$ [43]. To calculate $g^{i,\chi}_{\mathbf{k}}$, with $i=1\leavevmode\nobreak\ \&\leavevmode\nobreak\ 2$, we use Eq. 64. Keeping terms up-to linear order in electric ($\mathbf{E}$) field, the Boltzmann transport equation can be written as,
@@
$\left[\left(\frac{\partial g^{1,\chi}}{\partial\mathbf{k}}\right)\cdot\left(\mathbf{E}+\frac{e\mathbf{B}\cdot\mathbf{E}}{\hbar}\leavevmode\nobreak\ \bm{\Omega}^{\chi}_{\mathbf{k}}\right)\right]$
@@
In the above two equations (Eq. 7 and  8), we have ignored the Lorentz force term presented in the expression of $\mathbf{\dot{k^{\chi}}}$, which can be justified by comparing the magnitude. We fix electric field along increasing $x$-direction and magnetic field can be rotated in $xz$-plane. Therefore, $\mathbf{E}=E(0,0,1)$ and $\mathbf{B}=B(\cos{\gamma},0,\sin{\gamma})$, i.e., for $\gamma=\pi/2$ both fields are parallel. In this case, only $z$-component of $\mathbf{\Lambda}^{\chi}_{\mathbf{k}}$ will be relevant. Eq. 7 and  8 can be written to:
@@
$\frac{\partial}{\partial\epsilon^{\chi}_{\mathbf{k}}}\left(\frac{\partial f_{0}^{\chi}}{\partial\epsilon^{\chi}_{\mathbf{k}}}\leavevmode\nobreak\ \Lambda^{1,\chi}_{\mathbf{k}}\right)\left[{v^{\chi,z}_{\mathbf{k}}}+\frac{eB\sin{\gamma}}{\hbar}(\mathbf{v^{\chi}_{k}}\cdot\mathbf{\Omega}^{\chi}_{k})\right]$
@@
$\left[\frac{\partial}{\partial\epsilon^{\chi^{\prime}}_{\mathbf{k^{\prime}}}}\left(\frac{\partial f_{0}^{\chi^{\prime}}}{\partial\epsilon^{\chi^{\prime}}{{}_{\mathbf{k^{\prime}}}}}\leavevmode\nobreak\ \Lambda^{1,\chi^{\prime}}_{\mathbf{k^{\prime}}}\right)\leavevmode\nobreak\ \Lambda^{2,\chi^{\prime}}_{\mathbf{k}^{\prime}}-\frac{\partial}{\partial\epsilon^{\chi}_{\mathbf{k}}}\left(\frac{\partial f_{0}^{\chi}}{\partial\epsilon^{\chi}_{\mathbf{k}}}\leavevmode\nobreak\ \Lambda^{1,\chi}_{\mathbf{k}}\right)\leavevmode\nobreak\ \Lambda^{2,\chi}_{\mathbf{k}}\right].$ (10)
@@ 3 $\sigma_{z}$ -1 -1 +1
$\frac{1}{\tau^{\chi}_{\mu}(\theta,\phi)}=\mathcal{V}\sum_{\chi^{\prime}}\Pi^{\chi\chi^{\prime}}\iint\frac{(k^{\prime})^{3}\sin{\theta^{\prime}}}{|\mathbf{v}^{\chi^{\prime}}_{k^{\prime}}\cdot{\mathbf{k^{\prime}}^{\chi^{\prime}}}|}d\theta^{\prime}d\phi^{\prime}\leavevmode\nobreak\ \mathcal{G}^{\chi\chi^{\prime}}(D^{\chi^{\prime}}_{\mathbf{k^{\prime}}})^{-1}.$ (13)
$\mathcal{V}\sum_{\chi^{\prime}}\Pi^{\chi\chi^{\prime}}\iint f^{\chi^{\prime}}(\theta^{\prime},\phi^{\prime})\leavevmode\nobreak\ \mathcal{G}^{\chi\chi^{\prime}}d\theta^{\prime}d\phi^{\prime}\times[d^{\chi^{\prime}}-h^{\chi^{\prime}}_{\mu}(\theta^{\prime},\phi^{\prime})\\ +a^{\chi^{\prime}}\cos\theta^{\prime}+b^{\chi^{\prime}}\sin\theta^{\prime}\cos{\phi^{\prime}}+c^{\chi^{\prime}}\sin{\theta^{\prime}}\cos{\phi^{\prime}}].$ (14)
@@
$d^{\chi}+a^{\chi}\cos{\phi}+b^{\chi}\sin{\theta}\cos{\phi}+c^{\chi}\sin{\theta}\sin{\phi}\\ =\sum_{\chi^{\prime}}\mathcal{V}\Pi^{\chi\chi^{\prime}}\iint f^{\chi^{\prime}}(\theta^{\prime},\phi^{\prime})\leavevmode\nobreak\ \mathcal{G}^{\chi\chi^{\prime}}d\theta^{\prime}d\phi^{\prime}\\ \times[d^{\chi^{\prime}}-h^{\chi^{\prime}}_{k^{\prime}}+a^{\chi^{\prime}}\cos{\theta^{\prime}}+b^{\chi^{\prime}}\sin{\theta^{\prime}}\cos{\phi^{\prime}}+c^{\chi^{\prime}}\sin{\theta^{\prime}}\sin{\phi^{\prime}}].\\$ (15)
@@
Here, $g^{\chi}_{\mathbf{k}}=f_{0}+g^{1,\chi}_{\mathbf{k}}+g^{2,\chi}_{\mathbf{k}}$. For eight unknowns ($d^{\pm 1},a^{\pm 1},b^{\pm 1},c^{\pm 1}$), equations LABEL:Boltzman_final and 16 are simultaneously solved with Eq. 13. Due to intricate structure of equations, all two-dimensional integrals with respect to $\theta^{\prime}$ and $\phi^{\prime}$, simultaneous equations’ solutions are carried out numerically. Returning back to the Eq. 10, expansion of each terms leads to:
$\mathcal{A}^{\chi}_{\mathbf{k}}(\theta,\phi)-\mathcal{B}^{\chi}_{\mathbf{k}}(\theta,\phi)\\ +\frac{\partial}{\partial\epsilon^{\chi}_{\mathbf{k}}}\left(\frac{\Lambda^{1,\chi}_{\mathbf{k}}}{\tau^{\chi}_{\mathbf{k}}(\theta,\phi)}\Lambda^{2,\chi}_{\mathbf{k}}\right)-\frac{1}{\tau^{\chi}_{\mathbf{k}}(\theta,\phi)}\frac{\partial\Lambda^{1,\chi}_{\mathbf{k}}}{\partial\epsilon^{\chi}_{\mathbf{k}}}\Lambda^{2,\chi}_{\mathbf{k}}\\ =\sum_{\chi^{\prime}}\left[\frac{\partial}{\partial\epsilon^{\chi}_{\mathbf{k}}}\left(\int\frac{d^{3}\mathbf{k^{\prime}}}{(2\pi)^{3}}\mathcal{D}^{\chi^{\prime}}_{\mathbf{k^{\prime}}}W^{\chi\chi^{\prime}}_{\mathbf{k}\mathbf{k^{\prime}}}\Lambda^{1,\chi^{\prime}}_{\mathbf{k^{\prime}}}\Lambda^{2,\chi^{\prime}}_{\mathbf{k^{\prime}}}\right)\right.\\ \left.-\int\frac{d^{3}\mathbf{k^{\prime}}}{(2\pi)^{3}}\mathcal{D}^{\chi^{\prime}}_{\mathbf{k^{\prime}}}W^{\chi\chi^{\prime}}_{\mathbf{k}\mathbf{k^{\prime}}}\frac{\partial\Lambda^{1,\chi^{\prime}}_{\mathbf{k^{\prime}}}}{\partial\epsilon^{\chi^{\prime}}_{\mathbf{k^{\prime}}}}\Lambda^{2,\chi^{\prime}}_{\mathbf{k^{\prime}}}\right],$ (17)
@@
Eq. 9 can be solved for $\Lambda^{1,\chi}$, which can then be used to solve for $\Lambda^{2,\chi}$ in Eq. 17, and then the distribution function is evaluated using Eq. 2 and  3 along $f^{\chi}_{\mathbf{k}}=f_{0}+g^{1,\chi}_{\mathbf{k}}+g^{2,\chi}_{\mathbf{k}}+O(E^{3})$. Once the distribution function is evaluated, the current density can be evaluated as:
@@
To evaluate all the different responses, $\mathbf{J}^{\mathrm{CNLH}}$ is written as [58]:
@@
with, $\alpha,\beta,\gamma=\{x,y,z\}$. A comparison between Eq. 19 and Eq. 20 allows us to identify the individual components of the third-rank nonlinear conductivity tensor, $\sigma^{\chi}_{\alpha\beta\gamma}$. For an applied electric field $\mathbf{E}=E\hat{z}$, the anomalous velocity, given by $\mathbf{v}^{\chi}_{\mathrm{anom}}\sim\mathbf{E}\times\mathbf{\Omega}^{\chi}_{\mathbf{k}}$, lies within the $xy$-plane. As the magnetic field is rotated within the $xz$-plane, the measurable Hall response manifests along the $y$-direction, corresponding to the $\sigma_{yzz}$ component. The contribution of $\mathcal{O}(E^{2})$ to the chiral anomaly-induced nonlinear Hall current $\mathbf{J}^{\mathrm{CNLH}}$ arises through the following relations: $g^{1,\chi}_{\mathbf{k}}\sim\mathbf{E}\cdot\mathbf{\Lambda}^{1,\chi}_{\mathbf{k}}$, $g^{2,\chi}_{\mathbf{k}}\sim\frac{\partial}{\partial\epsilon^{\chi}_{\mathbf{k}}}\left(\frac{\partial f_{0}^{\chi}}{\partial\epsilon^{\chi}_{\mathbf{k}}}\,\mathbf{\Lambda}^{1,\chi}_{\mathbf{k}}\cdot\mathbf{E}\right)\,\mathbf{\Lambda}^{2,\chi}_{\mathbf{k}}\cdot\mathbf{E}$, and the anomalous velocity $\mathbf{v}^{\chi}_{a}\sim\mathbf{E}\times\mathbf{\Omega}^{\chi}_{\mathbf{k}}$. The vector $\mathbf{\Lambda}^{1,\chi}_{\mathbf{k}}$ can generally be expressed as $\mathbf{\Lambda}^{1,\chi}_{\mathbf{k}}=\varphi(\theta,\phi,k(B,\gamma))$, where $\varphi$ denotes an arbitrary function determined by the specific form of $h^{\chi}_{\mathbf{k}}$. Upon examining the product of $\mathbf{v}^{\chi}_{a}$ and the distribution function $g^{\chi}_{\mathbf{k}}$, it becomes evident that the leading $\mathcal{O}(E^{2})$ contribution to $\mathbf{J}^{\mathrm{CNLH}}$ stems solely from $g^{1,\chi}_{\mathbf{k}}$. To go beyond this order and obtain higher-order corrections in current($\mathbf{J}$), one would need to solve Eq. 10 to determine $\mathbf{\Lambda}^{2,\chi}_{\mathbf{k}}$, which we leave as an open problem for future investigation. It is important to note that the magnetic field is oriented in the $xz$-plane, resulting in a nonlinear Hall response along the $y$-direction, i.e., $\sigma_{yzz}$. While an additional Hall component—namely, the planar Hall conductivity $\sigma_{xzz}$—is also possible, it is not considered in the current study and is left for future work.
II.2 Spin-orbit coupled (SOC) noncentrosymmetric metals (NCMs) with weak Zeeman field
[…]
$+\hbar\vartheta(k_{x}t_{x}+k_{z}t_{z})\sigma_{0}.$ (22)
[…]
where, $t_{x}^{\prime}=\left(t_{x}+\frac{\hbar^{2}}{2m}\frac{2\mathcal{M}\cos{\gamma}}{(\hbar\vartheta)^{2}}\right)$, $t_{z}^{\prime}=\left(t_{z}+\frac{\hbar^{2}}{2m}\frac{2\mathcal{M}\sin{\gamma}}{(\hbar\vartheta)^{2}}\right)$, and the constant energy shift is given by $H_{0}=\left(\frac{\mathcal{M}}{\hbar\vartheta}\right)^{2}\frac{\hbar^{2}}{2m}+t_{z}\mathcal{M}\sin{\gamma}+t_{x}\mathcal{M}\cos{\gamma}$. Thus, under a coordinate transformation that shifts the origin, the Zeeman field effectively renormalizes the tilt components along the $x$- and $z$-directions as $t^{\mathrm{Z}}_{x}=\frac{\hbar^{2}}{2m}\frac{2\mathcal{M}\cos{\gamma}}{(\hbar\vartheta)^{2}}$ and $t^{\mathrm{Z}}_{z}=\frac{\hbar^{2}}{2m}\frac{2\mathcal{M}\sin{\gamma}}{(\hbar\vartheta)^{2}}$, respectively. Additionally, it introduces a tilt-dependent constant term to the Hamiltonian. This constant shift alters the Fermi surface profile as a function of the angle $\gamma$. Equation 23 retains the structural form of the Hamiltonian used for SOC-NCMs in Ref. [35], allowing for straightforward diagonalization. The resulting energy eigenvalues are given by:
$+\hbar\vartheta(q_{x}t_{x}^{\prime}+q_{z}t_{z}^{\prime})+H_{\mathrm{0}},$ (24)
where $\chi=\pm 1$ denotes the two spin-orbit split bands. To obtain a constant energy contour—analogous to the case of Weyl semimetals—one must invert Eq. 24 with respect to $q$. In the transformed coordinate system, the expression for the orbital magnetic moment (OMM) retains a form similar to that of Weyl semimetals in the original coordinates. The inclusion of the OMM term renders Eq. 24 a cubic equation in $q$, yielding three possible solutions for each value of $\chi$. Among these, the specific solution for $q$ selected for analysis is plotted as a function of $\theta$ at $\phi=0$ in Fig. 4 for both $\chi=\pm 1$. The Fermi surface undergoes a chirality-dependent energy shift due to the OMM, i.e., $\epsilon^{\chi}_{\mathbf{k}}\rightarrow\epsilon_{\mathbf{k}}-\mathbf{m}^{\chi}_{\mathbf{k}}\cdot\mathbf{B}$. As a consequence of the modified dispersion relation, the velocity components are also altered and are given by the following expressions:
$v^{\chi}_{y}$ $=\left(\frac{\hbar}{m}\right)k_{y}+\vartheta\frac{k_{y}}{k}$
with $u^{\chi}_{2}=\frac{\chi e\vartheta B}{2\hbar}$.
@@ II.2 Spin-orbit coupled (SOC) noncentrosymmetric metals (NCMs) with weak Zeeman field
II.2.1 Changes in Maxwell-Boltzmann transport (MBT) theory due to shift of origin

### v3

@@ Appendix A Maxwell Boltzmann Transport formalism
where $f^{\chi}_{\mathbf{k}}=f_{0}+g^{1,\chi}_{\mathbf{k}}+g^{2,\chi}_{\mathbf{k}}+O(E^{3})$. $f_{0}$ is the standard Fermi-Dirac distribution, $g^{1,\chi}_{\mathbf{k}}$ and $g^{2,\chi}_{\mathbf{k}}$ are deviation up to the first order and second order in electric field $E$, respectively, which can be written in following form [PhysRevB.103.045105, PhysRevB.104.205124, PhysRevB.105.125131]:
@@
Here, ‘$n$’ is the impurity concentration, $\mathcal{V}$’ is the system volume, $u^{\chi}(\mathbf{k})$ is the spinor wave function, $U^{\chi\chi^{\prime}}_{\mathbf{k}\mathbf{k}^{\prime}}$ is the scattering potential profile, and $\epsilon_{F}$ is the Fermi energy. We choose $U^{\chi\chi^{\prime}}_{\mathbf{k}\mathbf{k}^{\prime}}$ such that it can include both magnetic and non-magnetic point-like scattering centers. In general, $U^{\chi\chi^{\prime}}_{\mathbf{k}\mathbf{k}^{\prime}}=U^{\chi\chi^{\prime}}\sigma_{i}$ with $i=0,x,y,z$, where $U^{\chi\chi^{\prime}}$ distinguishes the interband ($\chi\neq\chi^{\prime}$) and intraband ($\chi=\chi^{\prime}$) scattering. Here we would like to highlight that in our formalism, the relative strength of these two scattering channels can be tuned using the parameter defined as:
$\alpha=\frac{U^{\chi\chi^{\prime}\neq\chi}}{U^{\chi\chi^{\prime}=\chi}}.$ (12)
In this work, we consider the geometry represented in Fig. 12(a), where the electric field is fixed along the $z$-direction, and the magnetic field is rotated in the $xz$-plane, forming an angle $\gamma$ with the $x$-axis. Including the topological nature of the Hamiltonian, in the presence of electric ($\mathbf{E}$) and magnetic ($\mathbf{B}$) fields, semiclassical dynamics of the Bloch electrons are modified and governed by the following equation [son2012berry, knoll2020negative]:
@@
where, $\mathbf{p}=\hbar\mathbf{k}$, $\mathbf{v}_{\mathbf{k}}^{\chi}=\frac{1}{\hbar}\frac{\partial\epsilon^{\chi}(\mathbf{k})}{\partial\mathbf{k}}$ is band velocity, $\bm{\Omega}^{\chi}_{\mathbf{k}}=i\nabla_{\mathbf{k}}\times\langle u^{\chi}(\mathbf{k})|\nabla_{\mathbf{k}}|u^{\chi}(\mathbf{k})\rangle$ is the Berry curvature, and $\mathcal{D}^{\chi}=(1+e\mathbf{B}\cdot\bm{\Omega}^{\chi}/\hbar)^{-1}$ is factor by which density of states is modified due to presence of the Berry curvature. Self-rotation of the Bloch wave packet also gives rise to an orbital magnetic moment (OMM) $\mathbf{m}^{\chi}_{\mathbf{k}}=-\frac{ie}{2\hbar}\text{Im}\langle\nabla_{\mathbf{k}}u^{\chi}|[\epsilon_{0}(\mathbf{k})-\hat{H}^{\chi}(\mathbf{k})]|\nabla_{\mathbf{k}}u^{\chi}\rangle$ [xiao2010berry]. To calculate $g^{i,\chi}_{\mathbf{k}}$, with $i=1\penalty 10000\ \&\penalty 10000\ 2$, we use Eq. 13, 10. Keeping terms up-to linear order in electric ($\mathbf{E}$) field, the Boltzmann transport equation can be written as,
@@
$\left[\left(\frac{\partial g^{1,\chi}}{\partial\mathbf{k}}\right)\cdot\left(\mathbf{E}+\frac{e\mathbf{B}\cdot\mathbf{E}}{\hbar}\penalty 10000\ \bm{\Omega}^{\chi}_{\mathbf{k}}\right)\right]$
@@
The present study primarily aims to examine the influence of the chiral anomaly term, while the Lorentz force contribution is neglected. This is because we work in the weak-field, weak-Lorentz-force regime where the small parameter is the cyclotron factor given by
$\omega_{c}\tau\equiv\frac{eB}{m}\tau\ll 1$ (16)
where $\tau$ is the relevant transport time. In this regime, the Lorentz force term in $\mathbf{\dot{p}}$ is suppressed by $\omega_{c}\tau$ compared to the electric field term. However, the chiral anomaly term $(\mathbf{E\cdot B)\Omega_{k}^{\chi}}$ is not suppressed in the same way. For typical SOC-NCM parameters, one has $\omega_{c}\tau\ll 1$ in the metallic regime where the semi-classical description is valid. A large body of work is found in literature in the above mentioned metallic regime [ahmad2021longitudinal, sharma2016nernst, PhysRevB.107.115161, nandy2021chiral, li2021nonlinear, zeng2022chiral]. We fix electric field along increasing $z$-direction and magnetic field can be rotated in $xz$-plane. Therefore, $\mathbf{E}=E(0,0,1)$ and $\mathbf{B}=B(\cos{\gamma},0,\sin{\gamma})$, i.e., for $\gamma=\pi/2$ both fields are parallel. In this case, only $z$-component of $\mathbf{\Lambda}^{\chi}_{\mathbf{k}}$ will be relevant. Eq. 14 and 15 can be written to:
@@
$\frac{\partial}{\partial\epsilon^{\chi}_{\mathbf{k}}}\left(\frac{\partial f_{0}^{\chi}}{\partial\epsilon^{\chi}_{\mathbf{k}}}\penalty 10000\ \Lambda^{1,\chi}_{\mathbf{k}}\right)\left[{v^{\chi,z}_{\mathbf{k}}}+\frac{eB\sin{\gamma}}{\hbar}(\mathbf{v^{\chi}_{k}}\cdot\mathbf{\Omega}^{\chi}_{k})\right]$
@@
$\left[\frac{\partial}{\partial\epsilon^{\chi^{\prime}}_{\mathbf{k^{\prime}}}}\left(\frac{\partial f_{0}^{\chi^{\prime}}}{\partial\epsilon^{\chi^{\prime}}{{}_{\mathbf{k^{\prime}}}}}\penalty 10000\ \Lambda^{1,\chi^{\prime}}_{\mathbf{k^{\prime}}}\right)\penalty 10000\ \Lambda^{2,\chi^{\prime}}_{\mathbf{k}^{\prime}}-\frac{\partial}{\partial\epsilon^{\chi}_{\mathbf{k}}}\left(\frac{\partial f_{0}^{\chi}}{\partial\epsilon^{\chi}_{\mathbf{k}}}\penalty 10000\ \Lambda^{1,\chi}_{\mathbf{k}}\right)\penalty 10000\ \Lambda^{2,\chi}_{\mathbf{k}}\right].$ (18)
@@ 3 $\sigma_{z}$ -1 -1 +1
$\frac{1}{\tau^{\chi}_{\mu}(\theta,\phi)}=\mathcal{V}\sum_{\chi^{\prime}}\Pi^{\chi\chi^{\prime}}\iint\frac{(k^{\prime})^{3}\sin{\theta^{\prime}}}{|\mathbf{v}^{\chi^{\prime}}_{k^{\prime}}\cdot{\mathbf{k^{\prime}}^{\chi^{\prime}}}|}d\theta^{\prime}d\phi^{\prime}\penalty 10000\ \mathcal{G}^{\chi\chi^{\prime}}(D^{\chi^{\prime}}_{\mathbf{k^{\prime}}})^{-1}.$ (21)
$\mathcal{V}\sum_{\chi^{\prime}}\Pi^{\chi\chi^{\prime}}\iint f^{\chi^{\prime}}(\theta^{\prime},\phi^{\prime})\penalty 10000\ \mathcal{G}^{\chi\chi^{\prime}}\penalty 10000\ \Lambda^{1,\chi^{\prime}}_{\mu}/\tau^{\chi^{\prime}}_{\mu}\penalty 10000\ d\theta^{\prime}d\phi^{\prime}.$ (22)
@@
$d^{\chi}+a^{\chi}\cos{\phi}+b^{\chi}\sin{\theta}\cos{\phi}+c^{\chi}\sin{\theta}\sin{\phi}\\ =\sum_{\chi^{\prime}}\mathcal{V}\Pi^{\chi\chi^{\prime}}\iint f^{\chi^{\prime}}(\theta^{\prime},\phi^{\prime})\penalty 10000\ \mathcal{G}^{\chi\chi^{\prime}}d\theta^{\prime}d\phi^{\prime}\\ \times[d^{\chi^{\prime}}-h^{\chi^{\prime}}_{k^{\prime}}+a^{\chi^{\prime}}\cos{\theta^{\prime}}+b^{\chi^{\prime}}\sin{\theta^{\prime}}\cos{\phi^{\prime}}+c^{\chi^{\prime}}\sin{\theta^{\prime}}\sin{\phi^{\prime}}].\\$ (23)
@@
Here, $g^{\chi}_{\mathbf{k}}=f_{0}+g^{1,\chi}_{\mathbf{k}}+g^{2,\chi}_{\mathbf{k}}$. For eight unknowns ($d^{\pm 1},a^{\pm 1},b^{\pm 1},c^{\pm 1}$), equations LABEL:Boltzman_final and 24 are simultaneously solved with Eq. 21. Due to intricate structure of equations, all two-dimensional integrals with respect to $\theta^{\prime}$ and $\phi^{\prime}$, simultaneous equations’ solutions are carried out numerically. Returning back to the Eq. 18, expanding each term and using the properties of Dirac delta function and its derivative, we get the following equation:
$\mathcal{A}^{\chi}_{\mathbf{k}}(\theta,\phi)-\mathcal{B}^{\chi}_{\mathbf{k}}(\theta,\phi)\\ +\frac{\partial}{\partial\epsilon^{\chi}_{\mathbf{k}}}\left(\frac{\Lambda^{1,\chi}_{\mathbf{k}}}{\tau^{\chi}_{\mathbf{k}}(\theta,\phi)}\Lambda^{2,\chi}_{\mathbf{k}}\right)-\frac{1}{\tau^{\chi}_{\mathbf{k}}(\theta,\phi)}\frac{\partial\Lambda^{1,\chi}_{\mathbf{k}}}{\partial\epsilon^{\chi}_{\mathbf{k}}}\Lambda^{2,\chi}_{\mathbf{k}}\\ =\sum_{\chi^{\prime}}\left[\frac{\partial}{\partial\epsilon^{\chi}_{\mathbf{k}}}\left(\int\frac{d^{2}\mathbf{k^{\prime}}}{(2\pi)^{3}}\mathcal{D}^{\chi^{\prime}}_{\mathbf{k^{\prime}}}W^{\chi\chi^{\prime}}_{\mathbf{k}\mathbf{k^{\prime}}}\Lambda^{1,\chi^{\prime}}_{\mathbf{k^{\prime}}}\Lambda^{2,\chi^{\prime}}_{\mathbf{k^{\prime}}}\right)\right.\\ \left.-\int\frac{d^{2}\mathbf{k^{\prime}}}{(2\pi)^{3}}\mathcal{D}^{\chi^{\prime}}_{\mathbf{k^{\prime}}}W^{\chi\chi^{\prime}}_{\mathbf{k}\mathbf{k^{\prime}}}\frac{\partial\Lambda^{1,\chi^{\prime}}_{\mathbf{k^{\prime}}}}{\partial\epsilon^{\chi^{\prime}}_{\mathbf{k^{\prime}}}}\Lambda^{2,\chi^{\prime}}_{\mathbf{k^{\prime}}}\right],$ (25)
@@
Eq. 17 can be solved for $\Lambda^{1,\chi}$, which can then be used to solve for $\Lambda^{2,\chi}$ in Eq. 25, and then the distribution function is evaluated using Eq. 8 and 9 along $f^{\chi}_{\mathbf{k}}=f_{0}+g^{1,\chi}_{\mathbf{k}}+g^{2,\chi}_{\mathbf{k}}+O(E^{3})$. Once the distribution function is evaluated, the current density can be evaluated as:
@@
To evaluate all the different responses, $\mathbf{J}^{\mathrm{CNLH}}$ is written as [yao2024geometrical]:
@@
with, $\alpha,\beta,\gamma=\{x,y,z\}$. A comparison between Eq. 27 and Eq. 28 allows us to identify the individual components of the third-rank nonlinear conductivity tensor, $\sigma^{\chi}_{\alpha\beta\gamma}$. For an applied electric field $\mathbf{E}=E\hat{z}$, the anomalous velocity, given by $\mathbf{v}^{\chi}_{\mathrm{anom}}\sim\mathbf{E}\times\mathbf{\Omega}^{\chi}_{\mathbf{k}}$, lies within the $xy$-plane. As the magnetic field is rotated within the $xz$-plane, the measurable Hall response manifests along the $y$-direction, corresponding to the $\sigma_{yzz}$ component. The contribution of $\mathcal{O}(E^{2})$ to the chiral anomaly-induced nonlinear Hall current $\mathbf{J}^{\mathrm{CNLH}}$ arises through the following relations: $g^{1,\chi}_{\mathbf{k}}\sim\mathbf{E}\cdot\mathbf{\Lambda}^{1,\chi}_{\mathbf{k}}$, $g^{2,\chi}_{\mathbf{k}}\sim\frac{\partial}{\partial\epsilon^{\chi}_{\mathbf{k}}}\left(\frac{\partial f_{0}^{\chi}}{\partial\epsilon^{\chi}_{\mathbf{k}}}\,\mathbf{\Lambda}^{1,\chi}_{\mathbf{k}}\cdot\mathbf{E}\right)\,\mathbf{\Lambda}^{2,\chi}_{\mathbf{k}}\cdot\mathbf{E}$, and the anomalous velocity $\mathbf{v}^{\chi}_{a}\sim\mathbf{E}\times\mathbf{\Omega}^{\chi}_{\mathbf{k}}$. The vector $\mathbf{\Lambda}^{1,\chi}_{\mathbf{k}}$ can generally be expressed as $\mathbf{\Lambda}^{1,\chi}_{\mathbf{k}}=\varphi(\theta,\phi,k(B,\gamma))$, where $\varphi$ denotes an arbitrary function determined by the specific form of $h^{\chi}_{\mathbf{k}}$. Upon examining the product of $\mathbf{v}^{\chi}_{a}$ and the distribution function $g^{1,\chi}_{\mathbf{k}}$ in Eq. 27, it becomes evident that the leading $\mathcal{O}(E^{2})$ contribution to $\mathbf{J}^{\mathrm{CNLH}}$ stems from $g^{1,\chi}_{\mathbf{k}}$ in the direction perpendicular to the electric field $\mathbf{E}$. Another Hall contribution $\sim E^{2}$ is also possible due to the multiplication of the velocity term $\mathbf{v_{k}^{\chi}}$ and second-order distribution function in $E$ (Eq. 26). However, it requires solving Eq. 18 to determine $\mathbf{\Lambda}^{2,\chi}_{\mathbf{k}}$, which we leave as an open problem for future investigation. It is important to note that the magnetic field is oriented in the $xz$-plane, resulting in a nonlinear Hall response along the $y$-direction, i.e., $\sigma_{yzz}$. While an additional non-linear Hall component—namely, the planar Hall conductivity $\sigma_{xzz}$ and a longitudinal component $\sigma_{zzz}$ (both of which require $\mathbf{\Lambda}^{2,\chi}_{\mathbf{k}}$) is also possible, it is not considered in the current study and is left for future work.
@@
Appendix B Changes in Maxwell-Boltzmann transport (MBT) theory due to shift of origin